OpenStudy (anonymous):

Help Use the functions f(x) = 5x − 12 and g(x) = 11x + 2 to complete the function operations listed below. Part A: Find (f + g)(x). Show your work. (3 points) Part B: Find (f ⋅ g)(x). Show your work. (3 points) Part C: Find f[g(x)]. Show your work. (4 points)

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@DaBest21

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

Part A ad like terms:- 5x - 12 + 11x + 2 bring like terms together 5x + 11x - 12 + 2 = ?

OpenStudy (welshfella):

can you continue /

OpenStudy (anonymous):

I got 16x-10 as the answer, is that right or am i off?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

So what about B and C? How do i solve those

OpenStudy (welshfella):

Part A is correct

OpenStudy (welshfella):

Part B multiply;- (5x - 12)(11x + 2)

OpenStudy (anonymous):

55x2−122x−24 ?

OpenStudy (welshfella):

first multipky the second bracket by 5x then multiply by -12 i;ll start you off: (5x - 12)(11x + 2) = 5x(11x + 2) - 12(11x + 2) = 55x^2 + 10x - ?

OpenStudy (welshfella):

yes you are correct

OpenStudy (welshfella):

Part C ?

OpenStudy (anonymous):

Yes i need help with C

OpenStudy (welshfella):

for part C you replace the x in f(x) by g(x)

OpenStudy (welshfella):